OpenStudy (anonymous):

The height (in feet) of a rocket from ground level is given by the function f(t) = -16t2 + 160t. What is the instantaneous velocity of the rocket 3 seconds after it is launched?

OpenStudy (anonymous):

32 feet/second 64 feet/second 12 feet/second 10 feet/second 8 feet/second

OpenStudy (anonymous):

Is it 64 feet /second

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute the first drivative of f(t), what do you get?

OpenStudy (anonymous):

f'(t) = -32t + 160

OpenStudy (michele_laino):

ok! Now the requested velocity is given by the subsequent quantity: \[v = f'\left( 3 \right) - 32 \times 3 + 160 = ...?\]

OpenStudy (michele_laino):

oops.. \[v = f'\left( 3 \right) = - 32 \times 3 + 160 = ...?\]

OpenStudy (anonymous):

64

OpenStudy (michele_laino):

that's right!